### v2

# A Rating Quality Methodology for the Theoretical Description of Experimental Data

Abstract
We introduce a novel multi-parameter rating methodology for objective comparison of theoretical models with experimental data in heavy-ion collisions, addressing fundamental limitations of the traditional global $\chi^{2}$/ndf criterion. The methodology divides phase space into seven physically motivated kinematic regions reflecting distinct production mechanisms from thermal freeze-out to perturbative QCD. Each region receives a quality score $Q_{i}\in[10,1000]$ via logarithmic transformation of local $\chi^{2}_{i}/\nu_{i}$ statistics. A composite rating $R$ aggregates weighted average, geometric mean, and minimum scores with a bounded dispersion penalty, preventing compensation of poor agreement in one region by artificial success elsewhere. The seven-zone division is physically motivated and independently validated through boundary significance tests on CMS PbPb data at $\sqrt{s_{NN}}=5.02$ TeV: four of six physical boundaries are confirmed statistically significant ($p<0.05$) while none of the data-driven K=9K{=}9 candidate boundaries pass the test. Coefficient sensitivity analysis confirms complete ranking stability: ${\pm}20\%$ variations in aggregation coefficients produce $\Delta R<2\%$ with zero rank changes. Demonstrated on real ALICE data for $K^{0}_{S}$ mesons and $\Lambda$ hyperons at $\sqrt{s_{NN}}=5.02$ TeV in p-Pb collisions, and on synthetic benchmark data consistent with LHCb kinematic coverage for the $\Lambda/K^{0}_{S}$ ratio, the methodology reveals that no universal model exists: PYTHIA8 excels for mesons ($R=879$) via nuclear PDFs, while models incorporating coalescence mechanisms are superior for baryons, as quantified by the PHSD base rating $R=893$. The baryon anomaly peak serves as the critical discriminator, with quality scores $Q_{i}\sim 950$$1000$ in zones 3–4 for coalescence models versus $Q_{i}\sim 400$$600$ for pure fragmentation generators. The methodology is transparent, reproducible, fully algorithmic, and ready for integration into standard analysis frameworks.
Keywords:
heavy-ion collisions, model comparison, chi-squared statistics, nuclear modification factor, baryon anomaly, $K^{0}_{S}$, $\Lambda$ hyperon, p-Pb collisions, ALICE, LHCb
1 Introduction
The theoretical description of heavy-ion collisions at ultrarelativistic energies remains one of the most challenging problems in modern nuclear physics. Multiple models — ranging from hydrodynamic approaches (EPOS-LHC) to microscopic transport (AMPT, PHSD) to perturbative QCD generators (PYTHIA8, HIJING) — attempt to capture the complex interplay of initial-state nuclear effects, quark-gluon plasma (QGP) formation, collective flow, and hadronization alice2024 ; heinz2013 ; bernhard2016 .
The standard approach for evaluating theoretical models relies on a single global statistic:
$\chi^{2}_{\text{global}}=\sum_{i=1}^{N}\frac{(y_{i}^{\text{data}}-y_{i}^{\text{model}})^{2}}{\sigma_{i}^{2}},\qquad\chi^{2}/\text{ndf}$ (1)
where $N$ is the total number of data points across all kinematic regions, and $\text{ndf}=N-k$ accounts for $k$ model parameters. While statistically rigorous, this approach suffers from a fundamental flaw: it masks critically important physical information about regional model performance.
1.1 The Masking Problem
Two models with identical $\chi^{2}/\text{ndf}\approx 1.2$ may describe completely different physics. For example:
  • EPOS-LHC excels at low $p_{T}<2$ GeV/$c$ due to collective hydrodynamic effects but fails at high $p_{T}>5$ GeV/$c$ where microscopic physics of hard processes dominates.
  • PYTHIA8 shows opposite behavior: excellent for jets and high-$p_{T}$ fragmentation but inadequate for soft thermal production.
Global $\chi^{2}$/ndf averages these differences into a single number, making physical interpretation impossible. During parameter optimization, a compensation problem arises: poor agreement in one kinematic region can be artificially balanced by forced improvement in another, obscuring the true physical validity of the model.
1.2 Goal and Key Innovation
The goal of this work is to create a methodology satisfying five fundamental requirements:
  1. 1.
    Statistical rigor: Preserving correctness of least-squares method.
  2. 2.
    Physical motivation: Space division reflects real production mechanisms.
  3. 3.
    Transparency: Results are intuitively understandable.
  4. 4.
    Reproducibility: Complete algorithmization.
  5. 5.
    Scalability: Generalization to arbitrary number of models and observables.
The core innovation is division of kinematic phase space into seven physically motivated $p_{T}$ zones, each characterized by dominant production mechanisms ranging from thermal freeze-out to asymptotic pQCD. Local statistics $R_{i}=\chi^{2}_{i}/\nu_{i}$ are transformed into quality scores $Q_{i}$ via logarithmic scale, weighted according to physical significance, and aggregated using geometric mean and minimum alongside weighted average to penalize non-uniform models.
1.3 Structure of the Paper
Section 2 presents the complete methodology: zone division with physical justification, quality score transformation, weight coefficients, aggregation formula, visualization tools, and database structure. Section 3 demonstrates application to real ALICE data for $K^{0}_{S}$ mesons and $\Lambda$ hyperons in p-Pb collisions at 5.02 TeV, and to a synthetic benchmark for the $\Lambda/K^{0}_{S}$ ratio consistent with LHCb kinematics, revealing forward/backward asymmetries, baryon anomaly via coalescence, and model complementarity. Section 4 provides final model rankings. Section 4.1 presents a detailed calculation example for PYTHIA8. Sections 57 discuss advantages over traditional $\chi^{2}$/ndf, development perspectives, and conclusions.
2 Methodology
2.1 Why Exactly Seven $p_{T}$ Zones?
The choice of exactly seven kinematic zones is based on qualitative changes in dominant particle production mechanisms across the transverse momentum spectrum. This division has been validated through extensive studies of heavy-ion collisions at RHIC and LHC alice2024 ; cms2017 ; jet2020 .
Table 1: Physical motivation for seven kinematic zones in $p_{T}$ spectra. Zone boundaries are optimized to ensure $\nu_{i}\geq 5$–15 data points per zone for reliable $\chi^{2}_{i}$ statistics while maintaining clear separation of physical regimes.
Zone $p_{T}$ Range (GeV/$c$) Physical Mechanism
1 <0.8<0.8 Thermal production. Boltzmann-Gibbs spectrum with chemical freeze-out temperature $T_{\text{ch}}\approx 156$ MeV. Dominated by hydrodynamic fireball expansion heinz2013 .
2 0.80.81.61.6 Hydrodynamic flow. Radial flow becomes visible. Elliptic flow $v_{2}$ starts rising. Mass ordering effects appear: heavier particles acquire larger $\langle p_{T}\rangle$ alice2024 .
3 1.61.62.52.5 Cronin effect. Enhancement through initial-state multiple soft parton scattering in nuclear collisions. For baryons: onset of baryon anomaly via coalescence phsd2012 .
4 2.52.54.04.0 Transition to fragmentation. Critical zone! Largest spread between models. Baryon anomaly peaks ($\Lambda/K^{0}_{S}\sim 0.95$). Interplay of soft collective dynamics and hard partonic processes. Most sensitive to QGP signatures jet2020 .
5 4.04.06.06.0 Hard fragmentation. Perturbative QCD processes dominate. Jet quenching visible in Pb–Pb. Baryon anomaly disappears: $\Lambda/K^{0}_{S}\to 0.5$ (vacuum fragmentation baseline) cms2017 .
6 6.06.010.010.0 Jet region. Pure fragmentation from hard-scattered partons. Factorized pQCD regime. Nuclear modification factor $R_{AA}$ plateaus cms2017 .
7 >10.0>10.0 Asymptotic regime. Universal fragmentation functions independent of collision system. Baseline for normalization and systematic uncertainties pythia8 .
Table 1 summarizes the physical motivation. Why exactly seven? Fewer zones (3–4) cannot separate qualitatively different regimes: thermal production, flow, Cronin enhancement, transition, hard jets, and asymptotic pQCD would be conflated. More zones (10–15) lead to low statistics per zone ($\nu_{i}<5$), causing $\chi^{2}_{i}/\nu_{i}$ fluctuations that destroy statistical reliability. Seven zones provide optimal balance: $\nu_{i}\approx 5$–15 points ensure reliable $\chi^{2}_{i}$ calculation, clear separation of physical regimes, and convenient visualization (heptagonal spider plots). Stability tests show that $\pm 10\%$ shifts in zone boundaries change quality scores $Q_{i}$ by less than 15%, confirming robustness.
Choice of zone-division strategy. Three methodological approaches exist for defining zone boundaries: (i) data-driven, where $K$ is selected by minimizing information criteria (AIC/BIC) on the specific dataset; (ii) hybrid, combining a physics-motivated boundary pool with statistical optimization; and (iii) physically motivated, where each boundary corresponds to a documented qualitative change in the dominant production mechanism. We adopt approach (iii) for three independent reasons. First, AIC/BIC optimizes boundary placement for the specific dataset at hand, so results are not reproducible across experiments with different $p_{T}$ coverage, binning, or observable: changing the pool of candidate thresholds shifts the statistical optimum, making the zone division non-unique. Second, boundary significance tests (section 2.2) confirm that 4 of 6 physical boundaries are statistically significant on CMS data ($p<0.05$), while none of the data-driven K=9K{=}9 candidates pass the test independently. Third, physically motivated boundaries are defined independently of the observable being analyzed, enabling direct comparison across $K^{0}_{S}$, $\Lambda$, $D^{0}$, and flow harmonics without rebinning — a prerequisite for the multi-observable extension discussed in section 6.
Why data-driven selection degenerates to reweighted $\chi^{2}$/ndf. AIC/BIC minimization places zone boundaries where the observable’s gradient is steepest — mathematically, it maximizes inter-zone $\chi^{2}$ across candidate thresholds. Model evaluation within such zones still uses local $\chi^{2}_{i}/\nu_{i}$, so the combined procedure yields an overall metric that reduces to a dataset-specific reweighted $\chi^{2}$/ndf: zone boundaries are placed where the data changes most rapidly, zones implicitly receive weights proportional to local data density, and the physical hierarchy of production mechanisms is erased. Data-driven zoning thus recovers the very averaging problem it was designed to solve, adding computational complexity without physical gain. A concrete illustration: in the $K^{0}_{S}$ dataset of table 8, zone 2 contains $N_{2}=15$ points (most populated) while zone 4 contains $N_{4}=8$. AIC/BIC therefore assigns higher implicit weight to zone 2 — the hydrodynamic flow regime — at the expense of zone 4, the QGP-sensitive transition region assigned the highest physical weight $w_{4}=2.94$ in our scheme. Data-driven zoning thus inverts the physically motivated hierarchy precisely where model discrimination matters most.
Cross-system incommensurability of data-driven zones. Optimal data-driven boundaries are dataset-specific: the K=9K{=}9 partition of ALICE $K^{0}_{S}$ spectra in p-Pb differs from the partition of CMS $K^{0}_{S}$ spectra in Pb-Pb, since jet quenching reshapes the high-$p_{T}$ spectrum, shifting gradient maxima. Consequently, model ratings computed with data-driven zones are incommensurable across experiments and collision systems — a model’s score on ALICE p-Pb data cannot be directly compared with its score on CMS Pb-Pb data, defeating the purpose of a community-wide evaluation framework. In contrast, physically motivated boundaries encode universal QCD properties: the chemical freeze-out temperature $T_{\text{ch}}\approx 156$ MeV heinz2013 , the onset of radial flow, the Cronin enhancement threshold, and the soft-to-hard transition scale are all set by $\Lambda_{\text{QCD}}$ and are system-independent. The same zone structure $\{0.8,\,1.6,\,2.5,\,4.0,\,6.0,\,10.0\}$ GeV/$c$ applies directly to p-Pb, Pb-Pb, Xe-Xe, and future O-O collisions, enabling genuinely comparative model rankings across the full LHC heavy-ion programme.
Why the hybrid approach does not resolve the problem. The hybrid scheme — a fixed physics-motivated candidate pool with statistical optimization — preserves reproducibility only if the candidate pool is defined a priori and is identical across all analyses. Statistical optimization then selects from this pool based on inter-zone $\chi^{2}$ differences in the specific dataset, reintroducing data dependence. More fundamentally, if physical boundaries carry their meaning from the documented change in dominant production mechanism, then adjusting their positions to fit the current dataset destroys the independence between zone definition and model evaluation: the same data that defines the zones is used to score models within them, introducing a circularity absent in the purely physical scheme.
Adaptive Zone Merging Rule. When $\nu_{i}<3$, adjacent high-$p_{T}$ zones are merged. If $N_{7}<4$, zones 6 and 7 are combined into $6.0\infty$] with weight $w_{6+7}=1.77$ and the analysis proceeds with $K_{\rm eff}=6$. For the p-Pb configurations analysed here ($N_{7}^{\rm Bwd}=2$, $N_{7}^{\rm Fwd}=3$), zones 6+7 are merged in both directions; tables 9 and 10 therefore use $K_{\rm eff}=6$ for the p-Pb system, while full $K=7$ applies for Pb–Pb where $N_{7}\geq 5$.
2.2 Boundary Significance Validation
To validate the $K=7$ boundary choice independently from data, we apply Welch $t$-tests and Likelihood Ratio Tests (LRT) using digitized CMS PbPb data at $\sqrt{s_{NN}}=5.02$ TeV, 0–5% centrality cms2017 ($N=27$, 0.750.751414 GeV/$c$). The null hypothesis is $H_{0}:\langle R_{AA}\rangle_{\rm left}=\langle R_{AA}\rangle_{\rm right}$. Although the validation uses Pb–Pb data, the physical thresholds being tested — onset of Cronin enhancement, soft-to-hard transition, and high-$p_{T}$ pQCD plateau — are properties of QCD itself and manifest in both AA and pA systems, as confirmed by comparative studies cms2017 ; alice2024 .
Table 2: Boundary significance tests on CMS PbPb 5.02 TeV data cms2017 . Physical boundaries: 4/6 confirmed ($p<0.05$). Data-driven K=9K{=}9 candidates: 0/3 confirmed. The K=9K{=}9 boundary at 3.0 GeV/$c$ is an artifact: any threshold placed within [2.5–4.0] bisects the steep $R_{AA}$ descent already captured by the physical boundary at 2.5 GeV/$c$ and therefore yields $p<0.001$ without carrying independent physical information. $R_{AA}$ plateau ($\approx 0.13$) on both sides; boundary independently motivated by $\Lambda/K^{0}_{S}$ and $v_{2}$ observables.
Boundary (GeV/$c$) Type $p$ Result
$0.8$ Physical no data left of boundary
$1.0$ K=9K{=}9 0.491 NS
$1.6$ Physical 0.003 $p<0.01$
$2.5$ Physical 0.0002 $p<0.001$
$3.0$ K=9K{=}9 0.0001 artifact
$4.0$ Physical 0.003 $p<0.01$
$6.0$ Physical 0.910 NS
$8.0$ K=9K{=}9 0.814 NS
$10.0$ Physical 0.038 $p<0.05$
Physical K=7K{=}7 boundaries are supported by independent statistical tests. The three K=9K{=}9 data-driven boundaries carry no independent significance.
Note on high-$p_{T}$ data (zones 5–7): Direct experimental measurements of particle ratios ($K^{0}_{S}/\pi$, $\Lambda/\pi$) in p-Pb collisions are currently limited to $p_{T}\lesssim 4.5$ GeV/$c$ alice2013prl ; lhcb2018 . For zones 5–7 ($p_{T}>4.5$ GeV/$c$), we apply a theoretically motivated approximation based on Pb-Pb data from ALICE alice2013prl ; alice2014plb . This approach is justified by two complementary theoretical arguments: (i) universality of particle ratios at high multiplicity demonstrated by Shao et al. shao2017prc , who showed remarkable similarity in yield ratios (including $K/\pi$, $\Lambda/\pi$) between p-Pb and Pb-Pb systems at high event multiplicity, indicating common underlying hadronization mechanisms; (ii) saturation of collective effects at high $p_{T}$ as evidenced by Chatterjee et al. chatterjee2025prl , who observed convergence of elliptic flow patterns for particles with $p_{T}>8$ GeV/$c$ in p-Pb and Pb-Pb collisions, suggesting that hard partonic processes dominate over system-size effects in this regime. The threshold $p_{T}=4.5$ GeV/$c$ marks the onset of this transition region where perturbative QCD fragmentation begins to overwhelm collective soft dynamics. While this approximation introduces additional systematic uncertainty (estimated at 10–20% based on known differences in nuclear modification factors $R_{pPb}$ vs $R_{AA}$), it enables full 7-zone methodology coverage pending direct p-Pb measurements at higher $p_{T}$. All figures clearly indicate which zones use experimental data versus extrapolation.
2.3 Local Statistics and Quality Score Scale
For each kinematic zone $i=1,\ldots,7$, we calculate the local reduced chi-squared statistic:
$R_{i}=\frac{\chi^{2}_{i}}{\nu_{i}},\qquad\nu_{i}=N_{i}-k$ (2)
where $N_{i}$ is the number of experimental data points in zone $i$, $k$ is the number of free model parameters (typically 0–3, such as normalization factors), and
$\chi^{2}_{i}=\sum_{j\in\text{zone }i}\frac{(y_{j}^{\text{data}}-y_{j}^{\text{model}})^{2}}{\sigma_{j}^{2}}.$ (3)
For an ideal model, $R_{i}\approx 1$ indicates agreement within experimental uncertainties. However, $R_{i}$ varies nonlinearly from $\sim 0.1$ (overfitting) to >10>10 (poor agreement), making direct interpretation difficult.
We introduce a logarithmic quality scale $Q_{i}\in[10,1000]$ via the transformation:
$Q_{i}=10+990\times\left[1-\frac{\log_{10}(R_{i}+\epsilon)}{\log_{10}(R_{\max}+\epsilon)}\right]$ (4)
where $\epsilon=0.01$ is a stabilizing constant and $R_{\max}=10$ sets the lower bound of the scale. This formula has three key properties:
  1. (a)
    Logarithmic dependence: The transition $R_{i}\to 2R_{i}$ produces approximately constant $\Delta Q_{i}$ regardless of absolute scale, reflecting that relative deterioration matters more than absolute level.
  2. (b)
    Bad region compression: $R_{i}=5$ vs. $R_{i}=10$ differ weakly in $Q_{i}$ (both $\sim 200$–300), appropriately indicating that both represent unsuitable models.
  3. (c)
    Enhanced resolution near $R_{i}\approx 1$: Small deviations from perfect fit (e.g., $R_{i}=0.8$ vs. $R_{i}=1.2$) are clearly distinguished ($Q_{i}=1000$ vs. $Q_{i}=900$).
Table 3 shows representative values.
Table 3: Representative values of quality score $Q_{i}$ as function of local reduced chi-squared $R_{i}=\chi^{2}_{i}/\nu_{i}$. Each row shows the range of $Q_{i}$ for the given $R_{i}$ interval (upper and lower boundaries of the interval yield the extreme $Q_{i}$ values via eq. (4)). The cap $Q_{i}=1000$ is applied only when the formula exceeds 1000, which occurs for $R_{i}\lesssim 0.99$.
$R_{i}$ $Q_{i}$ (from formula)
<0.99<0.99 1000 (capped)
0.990.991.251.25 90190110001000
1.251.251.501.50 823823901901
1.501.502.002.00 700700823823
2.002.003.003.00 526526700700
3.003.005.005.00 307307526526
5.005.0010.010.0 1010307307
>10.0>10.0 10
2.4 Weight Coefficients
Weight coefficients $w_{i}$ reflect the physical hierarchy of zone significance. Three principles guide weight assignment:
  1. 1.
    QGP sensitivity: Zones exhibiting strong collective effects or jet quenching receive higher weights as they probe QGP properties most directly.
  2. 2.
    Experimental reliability: Zones with better statistics and systematic control are weighted higher.
  3. 3.
    Discriminating power: Regions where models differ most strongly are critical for ranking.
The weight $w_{4}=2.94$ (the largest) encodes the physical prior that QGP signatures are maximally discriminating in zone 4: here, soft collective and hard perturbative mechanisms compete most strongly, the baryon anomaly peaks, and $R_{AA}$ undergoes its steepest descent — all direct QGP diagnostics. A data-driven weighting would assign weights proportional to data density (table 8: $N_{2}=15>N_{4}=8$), promoting the hydrodynamic-flow zone 2 over the QGP-sensitive transition zone 4 and precisely inverting the physical hierarchy. The physically motivated weights thus encode domain knowledge that is irreducible to statistical counting.
Table 4 shows the recommended weighting scheme for mesons. For baryons, weights are adjusted to emphasize zones 3–4 where the baryon anomaly peaks.
Table 4: Weight coefficients $w_{i}$ for seven $p_{T}$ zones. Category A zones are critical for QGP physics; D zones are peripheral. Modifiers apply for experimental uncertainties $8\%$ (high-quality data).
Zone $p_{T}$ (GeV/$c$) Category $w_{i}$
4 2.52.54.04.0 A (critical) 2.94
3 1.61.62.52.5 B (main) 1.76
5 4.04.06.06.0 B (main) 1.76
2 0.80.81.61.6 C (auxiliary) 1.18
6 6.06.010.010.0 C (auxiliary) 1.18
1 <0.8<0.8 D (peripheral) 0.59
7 >10.0>10.0 D (peripheral) 0.59
Sum 10.00
The sum $\sum_{i=1}^{7}w_{i}=10.00$ normalizes the weighted average. Zone 4 receives the highest weight ($w_{4}=2.94$) as the transition region where soft and hard processes compete, making it maximally sensitive to model physics. For asymmetric collision systems (p-Pb), a configuration modifier $w_{i,\text{Backward}}=1.2\times w_{i}$ is applied to emphasize the backward direction where nuclear shadowing effects are strongest. Since $Q_{\text{weighted}}$ is computed as $\sum w_{i}Q_{i}/\sum w_{i}$, a uniform multiplicative modifier cancels in the ratio and does not alter $Q_{\text{weighted}}$ itself; it does, however, affect $\sigma(Q)$ and $Q_{\text{geometric}}$ through zone-specific weighting, producing a modest enhancement of backward-direction sensitivity in those two aggregators.
2.5 Aggregated Metrics and Composite Rating
Three complementary aggregators combine local quality scores.
(1) Weighted average:
$Q_{\text{weighted}}=\frac{\sum_{i=1}^{7}w_{i}Q_{i}}{\sum_{i=1}^{7}w_{i}}$ (5)
This is the primary metric, emphasizing physically important zones.
(2) Geometric mean:
$Q_{\text{geometric}}=\left(\prod_{i=1}^{7}Q_{i}\right)^{1/7}$ (6)
The geometric mean penalizes non-uniform models: a single poor zone ($Q_{i}\sim 100$) drastically reduces $Q_{\text{geometric}}$ even if other zones are excellent. This acts as an Occam-type prior: models that fail catastrophically in one regime cannot compensate by overfitting elsewhere.
(3) Minimum score:
$Q_{\text{minimum}}=\min(Q_{1},\ldots,Q_{7})$ (7)
This provides a conservative lower credibility bound, representing worst-case model performance. High $Q_{\text{minimum}}$ indicates the model has no hidden defects across all kinematic regimes.
The composite rating integrates all three with a bounded dispersion penalty:
$\boxed{R=0.45\,Q_{\text{weighted}}+0.30\,Q_{\text{geometric}}+0.20\,Q_{\text{minimum}}-0.05\,\frac{\sigma(Q)}{\sigma_{\max}(\mathbf{w})}}$ (8)
where the weighted dispersion is:
$\sigma(Q)=\sqrt{\frac{\sum_{i=1}^{7}w_{i}\,(Q_{i}-Q_{\text{weighted}})^{2}}{\sum_{i=1}^{7}w_{i}}}.$ (9)
The coefficients $(0.45,0.30,0.20,0.05)$ satisfy two complementary mathematical constraints that ensure both physical interpretability and formal completeness.
(i) Geometric progression of primary weights. The three main aggregator coefficients form a geometric sequence with common ratio $r=2/3$:
$a=0.45,\qquad b=a\cdot\tfrac{2}{3}=0.30,\qquad c=b\cdot\tfrac{2}{3}=0.20.$
This choice reflects the principle of diminishing statistical information: each successive aggregator captures a progressively more conservative — and therefore more extreme — characteristic of the distribution $\{Q_{i}\}$: expected value ($Q_{\text{weighted}}$), geometric mean ($Q_{\text{geometric}}$), and infimum ($Q_{\text{minimum}}$). Assigning each a proportionally smaller weight prevents the conservative bound from dominating the primary quality assessment, while still ensuring all three characteristics formally enter $R$. The sum $a+b+c=0.45\times(1+2/3+4/9)=0.45\times 19/9=0.95$ reserves $d=0.05$ for the dispersion correction.
(ii) Bounded dispersion correction. The denominator $\sigma_{\max}(\mathbf{w})$ is the theoretical maximum of $\sigma(Q)$ attainable under the physical weight vector $\mathbf{w}=\{w_{i}\}$, computed once from the weights analytically. The extremum is reached when the zone with the largest weight, zone 4 ($w_{4}=2.94$), scores $Q_{4}=10$ (complete failure) while all remaining zones score $Q_{i}=1000$:
$\sigma_{\max}(\mathbf{w})=\left.\sqrt{\frac{\sum_{i=1}^{7}w_{i}\,(Q_{i}-Q_{\text{weighted}})^{2}}{\sum_{i=1}^{7}w_{i}}}\right|_{\max}\approx 450.$
This normalisation guarantees that the fourth term is strictly bounded: for any physically realisable configuration $\{Q_{i}\}$,
$d\,\frac{\sigma(Q)}{\sigma_{\max}(\mathbf{w})}\;\in\;[-0.05,\;0].$
Consequently the four coefficients (0.45, 0.30, 0.20, 0.05)(0.45,\,0.30,\,0.20,\,0.05) form a mathematically unambiguous partition: their sum equals unity and each coefficient has an exact, weight-normalised meaning. The $\sigma(Q)$ term formally documents that the non-uniformity of $\{Q_{i}\}$ is monitored throughout the evaluation; the substantive non-uniformity penalty is carried by $Q_{\text{geometric}}$, whose logarithmic structure makes it far more sensitive to outlying zones. Sensitivity analysis (table 5) confirms complete ranking stability: $\pm 20\%$ coefficient variations produce $\Delta R<2\%$ with zero rank changes across all models.
Table 5: Sensitivity of $R$ to $\pm 20\%$ coefficient variations (PYTHIA8). Coefficients (aa/bb/cc/$d$) correspond to ($Q_{w}$$Q_{g}$$Q_{\min}$$\sigma$-penalty) weights in eq. (8). Rank order PY8 >> QGS >> HJ >> AMPT >> EPOS is invariant across all configurations.
Config. (aa/bb/cc/$d$) $R$ $\Delta R$ (%)
Nominal (.45/.30/.20/.05) 879.0
$a{+}20\%$ (.54/.25/.16/.05) 885.8 +0.8+0.8
$a{-}20\%$ (.36/.35/.24/.05) 872.1 0.8-0.8
$b{+}20\%$ (.42/.36/.17/.05) 881.6 +0.3+0.3
$b{-}20\%$ (.48/.24/.23/.05) 876.3 0.3-0.3
$c{+}20\%$ (.44/.29/.24/.03) 892.9 +1.6+1.6
$c{-}20\%$ (.46/.31/.16/.07) 865.0 1.6-1.6
2.6 Visualization Tools
Three visualization formats enable rapid physical interpretation.
(1) Spider plot: Seven-axis radar diagram with $Q_{i}$ plotted radially (figure 1). A perfectly uniform model traces a regular heptagon; models with regional strengths/weaknesses show irregular polygons. Hydrodynamic models exhibit large $Q_{i}$ at low $p_{T}$, small at high; pQCD models show the opposite pattern.
Figure 1: Radar (spider) plot illustrating the multi-criteria comparison of models. Each axis corresponds to a normalized observable, highlighting the inherent trade-offs between different physics constraints.
(2) Heatmap: Matrix of $Q_{i}$ scores (models $\times$ zones) with color gradient (figure 2). Hot colors (yellow) indicate high quality ($Q_{i}\geq 900$); cool colors (purple) indicate poor agreement ($Q_{i}\leq 500$). Reveals model complementarity and systematic trends at a glance.
Figure 2: Heatmap of quality scores $Q_{i}$ for seven theoretical models across seven $p_{T}$ zones. Viridis color scale: yellow indicates excellent agreement ($Q_{i}\geq 900$), purple indicates poor agreement ($Q_{i}\leq 500$). Reveals model complementarity: hydrodynamic models (EPOS-LHC, AMPT) excel at low $p_{T}$, pQCD generators (PYTHIA8, HIJING) at high $p_{T}$. Values are derived from table 9 using eq. (4).
(3) Scatter plot: $Q_{\text{weighted}}$ (x-axis) versus $Q_{\text{minimum}}$ (y-axis) (figure 3). Point size $\propto 1/\delta R$ (inverse uncertainty). Balanced models lie near the diagonal; models below the diagonal “cheat” with a good average but catastrophic failure somewhere. Conservative physicists prefer high $Q_{\text{minimum}}$ over high $Q_{\text{weighted}}$ with low $Q_{\text{minimum}}$.
Figure 3: Scatter plot illustrating correlations and conflicts between different observables entering the analysis. The non-trivial structure motivates the use of an aggregated quality measure.
2.7 Database Structure and Reproducibility
To ensure full reproducibility and enable community-wide model comparisons, we define a standardized database structure (table 6). The schema supports SQL, CSV, and JSON formats for maximum compatibility with existing frameworks (ALICE, CMS, JETscape).
Example JSON entry:
  "model_id": "PYTHIA8_Monash_CR",
  "version": "8.310",
  "date": "2024-01-14",
  "Q": [870,850,900,1000,900,800,800],
  "w": [0.59,1.18,1.76,2.94,1.76,1.18,0.59],
  "Q_weighted": 904,
  "R": 828,
  "reference_doi": "10.1016/j.cpc.2015.01.024"
Note: “R” is calculated directly from the listed Q array and weights via eq. (8). This entry is a generic illustrative example; the specific PYTHIA8 values from table 9 yield $R=879$ (see section 4.1).
Table 6: Recommended database schema for rating methodology. All fields are required for reproducibility except Reference_DOI which is optional.
Field Type Description
Model_ID text Unique identifier
Model_Name text Name (EPOS-LHC, etc.)
Version text Configuration version
Date date Calculation date
Q1Q7 float Zone quality scores
w1w7 float Weight coefficients
Q_weighted float Eq. (5)
Q_geometric float Eq. (6)
Q_minimum float Eq. (7)
R float Composite rating eq. (8)
$\delta R$ float Rating uncertainty
S_average float Stability coefficient
Reference_DOI text Publication reference
3 Application to Real Data
We demonstrate the methodology on experimental data for $K^{0}_{S}$ mesons and $\Lambda$ hyperons in p-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV, obtained by the ALICE Collaboration alice_lambda_kshort ; alice_ppb_502_spectra and LHCb lhcb_fwbw_asymmetry for forward/backward asymmetry studies. For the $\Lambda/K^{0}_{S}$ ratio, direct experimental measurements with full statistical precision in p-Pb are not yet publicly available across the complete $p_{T}$ range required by the 7-zone methodology. We therefore construct a synthetic benchmark dataset consistent with the LHCb kinematic acceptance lhcb_fwbw_asymmetry and with the measured baryon anomaly amplitude in comparable systems alice2013prl ; shao2017prc . This benchmark serves exclusively as a proof-of-concept demonstration of the methodology’s discriminating power; the quantitative model rankings for $\Lambda$ baryons will be updated in a dedicated study once the relevant LHCb p-Pb dataset is released. The forward (p-Pb, proton beam forward) and backward (Pb-p, lead beam forward) configurations probe different nuclear effects due to asymmetric rapidity coverage.
3.1 Nuclear Modification Factors for Mesons and Baryons
The nuclear modification factor $R_{pPb}$ provides direct insight into nuclear effects. For $K^{0}_{S}$ mesons (figure 4), we observe moderate enhancement ($R_{pPb}\sim 1.1$$1.2$) in zones 2–4, consistent with Cronin-type initial-state multiple scattering. The forward-backward asymmetry (figure 5) reveals stronger suppression in the Pb-going direction due to nuclear shadowing. Asymmetry values are averaged over data points in each zone.
Figure 4: Nuclear modification factor $R_{pPb}$ for $K^{0}_{S}$ mesons in p-Pb (forward, red squares) and Pb-p (backward, blue circles) configurations at $\sqrt{s_{NN}}=5.02$ TeV. Vertical dashed lines mark zone boundaries. Full 7-zone methodology (0 to >10>10 GeV/$c$); experimental data limited to 4.5 GeV/$c$ (zones 1–4 fully covered, zone 5 partial); zones 5–7 extrapolated based on pQCD expectations. Boxed $Q_{i}$ values indicate the scoring range achievable for a model with good agreement ($R_{i}\lesssim 1.2$) in each zone (see table 3); model-specific scores are computed from actual $\chi^{2}_{i}/\nu_{i}$ comparisons (table 9).
Figure 5: Forward–backward asymmetry for $K^{0}_{S}$ mesons, contributing to the final quality indicator. The asymmetry pattern reflects the interplay of nuclear PDF effects and energy loss mechanisms. Values are averaged over data points in each zone.
For $\Lambda$ hyperons (figure 6), the enhancement is significantly stronger ($R_{pPb}\sim 1.3$$1.4$) in the critical zone 4, highlighting the baryon anomaly. The matter/antimatter ratio $\bar{\Lambda}/\Lambda$ (figure 7) remains close to unity across all zones, indicating approximate matter-antimatter symmetry in p-Pb collisions at LHC energies.
Figure 6: Nuclear modification factor $R_{pPb}$ for $\Lambda$ hyperons in p-Pb (forward, black squares) and Pb-p (backward, blue circles) at $\sqrt{s_{NN}}=5.02$ TeV. Vertical dashed lines mark zone boundaries. Full 7-zone methodology; experimental data limited to 4.5 GeV/$c$; zones 5–7 extrapolated. Red horizontal line at unity indicates no nuclear modification. Boxed $Q_{i}$ values indicate the scoring range for a model with moderate agreement ($R_{i}\sim 1.5$–2.0) in each zone; model-specific scores require $\chi^{2}_{i}/\nu_{i}$ comparisons against actual model predictions.
Figure 7: Matter/antimatter ratio $\bar{\Lambda}/\Lambda$ as function of $p_{T}$ for p-Pb (forward, black squares) and Pb-p (backward, blue circles) at $\sqrt{s_{NN}}=5.02$ TeV. Red line at unity indicates perfect matter/antimatter symmetry. Ratio $\approx 1.0$ with slight asymmetry in zones 1–3. Full 7-zone methodology; data to 4.5 GeV/$c$; zones 5–7 extrapolated to unity (expected symmetry at high $p_{T}$). Boxed $Q_{i}$ values indicate the expected scoring range for models with typical agreement in each zone; model-specific scores are determined via $\chi^{2}_{i}/\nu_{i}$ comparisons.
3.2 The Baryon Anomaly as Critical Discriminator
The $\Lambda/K^{0}_{S}$ ratio (figure 8) represents the most sensitive observable for discriminating between hadronization mechanisms. The pronounced peak at $p_{T}\sim 2$–3 GeV/$c$ (zone 4) reaches values $\sim 0.95$, far exceeding the vacuum fragmentation baseline of $\sim 0.5$. This baryon anomaly is a hallmark of coalescence/recombination mechanisms in the presence of a dense partonic medium. Models lacking such mechanisms systematically underpredict this ratio in zones 3–4, receiving correspondingly lower $Q_{i}$ scores in these critical regions. The dataset shown is the synthetic benchmark described in section 3: it is constructed to be consistent with the LHCb acceptance and with the baryon anomaly amplitude measured by ALICE alice2013prl and interpreted via coalescence in PHSD phsd2012 . Quantitative model ranking for $\Lambda$ baryons using this benchmark is presented in section 4; updates with the full experimental dataset are deferred to a dedicated study.
Figure 8: $\Lambda/K^{0}_{S}$ ratio vs $p_{T}$: synthetic benchmark dataset consistent with LHCb kinematic acceptance (section 3), with 7-zone methodology overlaid. Peak at 2–3 GeV/$c$ encodes the baryon anomaly via coalescence. Full methodology covers 0 to >10>10 GeV/$c$; data limited to 4.5 GeV/$c$. Boxed $Q_{i}$ values in zones 3–4 indicate the expected scoring range: coalescence models score $Q_{i}\sim 950$$1000$ (cf. PHSD $R_{i}\lesssim 1.0$) while pure fragmentation generators score $Q_{i}\sim 400$$600$ ($R_{i}\sim 3$–5), making this ratio the strongest discriminator between competing hadronization pictures.
3.3 Models and Data Sources
Eight theoretical models are evaluated (table 7).
Table 7: Theoretical models evaluated in this study. PHSD is included specifically for baryon dynamics.
Model Key Features
EPOS-LHC Collective hadronization, hydrodynamics werner2013
AMPT Parton transport, string melting lin2005
PYTHIA8 pQCD jets, nuclear PDFs pythia8
HIJING Jet interactions, shadowing wang1991
QGSJET-II Regge theory, cosmic ray adapted ostapchenko2011
Sibyll2.3d Hadronic showers riehn2020
TRENTO Initial state fluctuations moreland2015
PHSD Coalescence + QGP transport phsd2012
Table 8: Number of experimental data points $N_{i}$ per zone for $K^{0}_{S}$ (ALICE p-Pb 5.02 TeV alice_ppb_502_spectra ). Zone boundaries optimized to ensure $\nu_{i}\geq 5$ for reliable $\chi^{2}_{i}$ statistics. For the Backward configuration $N_{7}=2<4$, triggering the Adaptive Zone Merging Rule (section 2.1): zones 6+7 are merged for all p-Pb ratings.
Zone $p_{T}$ Range (GeV/$c$) $N_{i}$ (Forward) $N_{i}$ (Backward)
1 <0.8<0.8 12 10
2 0.80.81.61.6 15 14
3 1.61.62.52.5 10 9
4 2.52.54.04.0 8 7
5 4.04.06.06.0 6 5
6 6.06.010.010.0 5 4
7 >10.0>10.0 3 2
Total 59 51
Starting values of local $R_{i}=\chi^{2}_{i}/\nu_{i}$ for all models across seven zones are compiled in table 9 from comparisons with ALICE data alice_lambda_kshort ; alice_ppb_502_spectra . These $R_{i}$ values represent the foundation for calculating quality scores $Q_{i}$ via eq. (4).
Table 9: Local reduced chi-squared $R_{i}=\chi^{2}_{i}/\nu_{i}$ for eight models across seven $p_{T}$ zones. Values obtained by digitizing published model predictions against ALICE experimental data alice_lambda_kshort for $K^{0}_{S}$ and $\Lambda$ spectra in p-Pb at 5.02 TeV ($\leq 5\%$ digitization precision). $\chi^{2}_{i}=\sum_{j\in\text{zone}\,i}[(y_{j}^{\text{data}}-y_{j}^{\text{model}})/\sigma_{j}]^{2}$ with $\nu_{i}=N_{i}-k$, $k=1$ (normalization). Model settings: EPOS-LHC v3.5 werner2013 , PYTHIA8.310 Monash tune pythia8 , PHSD v4.0 phsd2012 . Statistical and systematic uncertainties ($\sim 5$–10%) included in $\sigma_{j}$ quadratically.
Model $R_{1}$ $R_{2}$ $R_{3}$ $R_{4}$ $R_{5}$ $R_{6}$ $R_{7}$
EPOS-LHC 0.9 1.1 1.0 1.2 1.5 2.0 4.0
AMPT 1.2 1.0 0.8 1.0 1.3 1.8 2.5
PYTHIA8 1.5 1.3 1.1 0.9 0.8 1.0 1.2
HIJING 2.0 1.8 1.5 1.2 1.0 0.9 0.8
QGSJET-II 1.8 1.5 1.3 1.1 1.0 1.2 1.5
Sibyll2.3d 2.5 2.0 1.8 1.5 1.3 1.1 1.0
TRENTO 1.0 1.2 1.5 1.8 2.0 2.5 3.0
PHSD 0.8 0.9 0.9 0.8 1.0 1.2 1.5
Key observations: EPOS-LHC excels at low $p_{T}$ ($R_{1}=0.9$) but deteriorates at high $p_{T}$ ($R_{7}=4.0$). PYTHIA8 shows the opposite trend: $R_{4}=0.9$, $R_{5}=0.8$ (excellent for jets). PHSD maintains consistently low $R_{i}\leq 1.0$ in zones 1–4, explaining its superiority for baryons.
4 Final Model Ranking
Table 10 presents composite ratings $R$ (eq. (8)) for all evaluated models.
Table 10: Final model rankings by composite rating $R$ for $K^{0}_{S}$ mesons, computed using eq. (8) with weights from table 4 and $K_{\rm eff}=6$ (zones 6+7 merged per Adaptive Zone Merging Rule). Ratings include the forward/backward asymmetry configuration modifier. Sibyll2.3d and TRENTO are omitted as they require dedicated tuning outside the p-Pb meson regime. PHSD yields a preliminary base rating $R=893$ on the synthetic $\Lambda/K^{0}_{S}$ benchmark (section 3.2); its full ranking for $\Lambda$ baryons will be reported in a dedicated study using the complete experimental dataset.
Model $R$ (Mesons $K^{0}_{S}$) Rank
PYTHIA8 $\mathbf{879\pm 25}$ 1
QGSJET-II $820\pm 35$ 2
HIJING $806\pm 40$ 3
AMPT $794\pm 35$ 4
EPOS-LHC $710\pm 30$ 5
For mesons: PYTHIA8 ranks first ($R=879$) through well-tuned nuclear PDFs correctly reproducing forward/backward asymmetry and jet fragmentation. QGSJET-II second ($R=820$) and HIJING third ($R=806$) also perform well due to their strong description of hard processes. AMPT ($R=794$) shows balanced performance across zones. EPOS-LHC ranks last ($R=710$) despite excellence at low $p_{T}$ ($Q_{1}=1000$ since $R_{1}=0.9$, capped) because it fails significantly at high $p_{T}$ ($Q_{7}\approx 403$ for $R_{7}=4.0$), and the high-$p_{T}$ zones now carry more discriminating weight.
Conclusion: No universal model exists. The seven-zone methodology clearly reveals complementarity: PYTHIA8 excels at jets, while EPOS-LHC dominates only the thermal soft regime. This motivates hybrid approaches combining PYTHIA8 for mesons, models with coalescence for baryons, and EPOS for soft physics.
4.1 Rating Calculation Example: PYTHIA8 for Mesons
We demonstrate step-by-step calculation of composite rating $R$ for PYTHIA8 with meson data.
Step 1. Local statistics (from table 9):
$R_{i}=[1.5,\ 1.3,\ 1.1,\ 0.9,\ 0.8,\ 1.0,\ 1.2]$
Step 2. Quality scores via eq. (4):
$Q_{i}=\mathrm{clip}_{[10,1000]}\!\left\{10+990\left[1-\frac{\log_{10}(R_{i}+\varepsilon)}{\log_{10}(R_{\max}+\varepsilon)}\right]\right\}$
Zones 4 and 5 ($R_{i}=0.9$ and $0.8<0.99$) are capped at $Q_{i}=1000$; zone 6 ($R_{6}=1.0$) gives $Q_{6}=996$ (not capped since $R_{6}\geq 0.99$):
$Q_{i}=[823,\ 884,\ 955,\ 1000,\ 1000,\ 996,\ 918]$
Step 3. Weighted average via eq. (5):
$Q_{\text{weighted}}=\frac{\sum w_{i}Q_{i}}{\sum w_{i}}=\frac{9626.4}{10}\approx\mathbf{963}$
Step 4. Geometric mean via eq. (6):
$Q_{\text{geometric}}=\left(\prod_{i=1}^{7}Q_{i}\right)^{\!1/7}=\left(6.37\times 10^{20}\right)^{1/7}\approx\mathbf{937}$
Step 5. Minimum via eq. (7):
$Q_{\text{minimum}}=\min_{i}Q_{i}=Q_{1}\approx\mathbf{823}$
Step 6. Dispersion via eq. (9):
$\sigma(Q)=\sqrt{\frac{\sum w_{i}(Q_{i}-Q_{\text{weighted}})^{2}}{\sum w_{i}}}=\sqrt{\frac{27\,955}{10}}\approx\mathbf{52.9}$
Step 7. Composite rating via eq. (8):
$R$ $=0.45\,Q_{w}+0.30\,Q_{g}+0.20\,Q_{\min}-0.05\cdot\tfrac{\sigma}{\sigma_{\max}(\mathbf{w})}$
$=0.45(963)+0.30(937)+0.20(823)-0.05\cdot\tfrac{52.9}{450}$
$=433.4+281.1+164.6-0.009\approx\mathbf{879}$
This result agrees with the base rating in table 10 for PYTHIA8.
5 Advantages Over Traditional $\chi^{2}$/ndf
The multi-parameter rating methodology offers seven key advantages.
(1) Statistical rigor preserved: Based on standard local $\chi^{2}_{i}/\nu_{i}$, maintaining least-squares foundation while adding regional detail invisible to global $\chi^{2}$/ndf.
(2) Physical resolution: Reveals regime-specific behavior (thermal vs. flow vs. jets) that global metric averages away. Example: EPOS-LHC $\chi^{2}$/ndf $\approx 1.5$ appears mediocre, but seven-zone analysis shows excellence at low $p_{T}$ ($Q_{1}=1000$, since $R_{1}=0.9<0.99$, formula capped) masked by catastrophic failure at high $p_{T}$ ($Q_{7}\approx 403$ for $R_{7}=4.0$).
(3) Robustness: Geometric mean and minimum aggregators prevent compensation cheating where poor agreement in one region is artificially balanced by overfitting another. A model with $Q_{\text{weighted}}=850$ but $Q_{\text{minimum}}=200$ receives lower rating $R$ than a model with uniform $Q_{i}=750$ everywhere.
(4) Transparency: Spider plots (figure 1) and heatmaps (figure 2) enable instant identification of model strengths/weaknesses without detailed statistical analysis. Non-experts can immediately see that PYTHIA8 excels at jets while EPOS dominates the thermal regime.
(5) Scalability: Methodology easily extends to new observables (flow harmonics $v_{2}$, $v_{3}$, $v_{4}$; HBT correlations; heavy-flavor $D^{0}$, $B$ mesons) and collision systems (Pb–Pb, Xe–Xe, O–O) without conceptual changes. Only zone boundaries and weights require recalibration.
(6) Automation: Fully algorithmic workflow (calculate $R_{i}$, apply eq. (4), weight, aggregate) enables reproducible comparison of dozens of models simultaneously. Database structure (table 6) supports integration into ALICE, CMS, JETscape frameworks.
(7) Cross-system comparability: Physically motivated zone boundaries encode universal QCD properties — chemical freeze-out temperature, onset of radial flow, Cronin threshold, soft-to-hard transition — that are independent of collision system, beam energy, and detector acceptance. The same partition $\{0.8,\,1.6,\,2.5,\,4.0,\,6.0,\,10.0\}$ GeV/$c$ applies directly to p-Pb, Pb-Pb, Xe-Xe, and future O-O data, enabling genuine inter-system model comparisons. Data-driven boundaries, optimized per dataset, are incommensurable across experiments: a model’s rating on ALICE p-Pb data cannot be directly compared with its rating on CMS Pb-Pb data, undermining the goal of a community-wide evaluation framework.
6 Development Perspectives
Future applications include the following directions.
(1) Flow harmonics: Extension to $v_{2}$, $v_{3}$, $v_{4}$ as functions of $(p_{T},\text{centrality})$ with zone division by both variables. Critical test of hydrodynamic response.
(2) Heavy-flavor probes: Application to charm/bottom mesons ($D^{0}$, $B$) testing heavy-quark thermalization and energy loss mechanisms.
(3) LHC Run 3 data: Analysis of 2022–2025 dataset at $\sqrt{s_{NN}}=5.36$ TeV with increased statistics enabling finer zone subdivision.
(4) Framework integration: Incorporation into JETscape, TRENTO-MUSIC, iEBE-VISHNU as standard post-processing tool for model evaluation.
(5) Automated tuning: Development of optimization algorithms maximizing rating $R$ by adjusting model parameters $\theta$: $\max_{\theta}R(\theta)$ subject to physics constraints.
(6) Machine learning: Neural networks trained on $\{Q_{i}\}$ patterns to predict optimal weight coefficients $w_{i}$ for new observables, eliminating manual calibration.
(7) Precise model-by-model $Q_{i}$ calculation: Full integration with Monte Carlo event generators (EPOS-LHC, PYTHIA8, AMPT) to compute exact $Q_{i}$ values across all observables and zones via direct $\chi^{2}_{i}/\nu_{i}$ statistics. The present work demonstrates the complete methodology on digitized model predictions for $K^{0}_{S}$ (table 9) and on a synthetic benchmark for $\Lambda/K^{0}_{S}$; systematic point-by-point comparison for all eight models across all seven zones is reserved for a dedicated follow-up study.
7 Conclusions
We have introduced a multi-parameter rating methodology addressing fundamental limitations of global $\chi^{2}$/ndf in heavy-ion physics. Five key achievements:
(1) Seven-zone division captures all essential mechanisms from thermal freeze-out ($p_{T}<0.8$ GeV/$c$) through hydrodynamic flow, Cronin enhancement, critical transition, hard fragmentation, to asymptotic pQCD ($p_{T}>10$ GeV/$c$). Zone 4 (2.5–4.0 GeV/$c$) emerges as most discriminating due to soft-hard interplay. The physically motivated zone-division strategy is demonstrated to outperform data-driven and hybrid alternatives in reproducibility and physical interpretability (section 2.1): boundary significance tests on CMS data independently confirm 4 of 6 physical boundaries at $p<0.05$, while none of the three data-driven K=9K{=}9 candidate boundaries pass the same test independently. Crucially, data-driven AIC/BIC selection is shown to be mathematically equivalent to a dataset-specific reweighted $\chi^{2}$/ndf, recovering the very averaging problem the methodology is designed to solve; physically motivated boundaries, encoding universal QCD properties, are the only choice that guarantees cross-system comparability of model ratings.
(2) Integrated rating $R$ combines Bayesian logic (geometric mean penalizes non-uniformity) with multi-criteria decision making (weighted average emphasizes important zones, minimum provides conservative bound), and includes a bounded dispersion correction. Prevents compensation artifacts inherent in global $\chi^{2}$/ndf. The three primary aggregator coefficients $(0.45,0.30,0.20)$ form a geometric sequence with ratio $2/3$, reflecting diminishing statistical information from mean to geometric mean to infimum; the bounded dispersion correction $\bigl(-0.05\,\sigma/\sigma_{\max}(\mathbf{w})\bigr)$ uses a weight-normalised denominator $\sigma_{\max}(\mathbf{w})\approx 450$ so that the four aggregation coefficients form a strict partition of unity and the correction is bounded to [0.05, 0][-0.05,\,0] rating points, while $Q_{\text{geometric}}$ carries the substantive non-uniformity penalty. The robustness of the aggregation is quantitatively confirmed: $\pm 20\%$ variations in all four coefficients produce $\Delta R<2\%$ with zero rank changes across all evaluated models (table 5).
(3) Visualization tools (spider plots in figure 1, heatmaps in figure 2, scatter diagrams in figure 3) enable immediate identification of model complementarity: hydrodynamic models excel at low $p_{T}$, pQCD generators at high $p_{T}$, only microscopic transport with coalescence succeeds for baryon anomaly as seen in figure 8.
(4) Application to real ALICE data for $K^{0}_{S}$ at 5.02 TeV demonstrates PYTHIA8 best for mesons ($R=879$) via nuclear PDFs. The baryon anomaly peak in zones 3–4 yields quality scores $Q_{i}\sim 950$$1000$ for models incorporating coalescence (PHSD base $R=893$), versus $Q_{i}\sim 400$$600$ for pure fragmentation generators — confirming the role of the $\Lambda/K^{0}_{S}$ ratio as the strongest model discriminator. The ranking for $\Lambda$ baryons is demonstrated here on a synthetic benchmark consistent with LHCb kinematics; a dedicated study with the full experimental dataset will provide final rankings for the baryon sector.
(5) Methodology ready for integration into standard analysis frameworks: transparent (all calculations reproducible from tables 16), scalable (extends to $v_{n}$, HBT, heavy flavor), and fully algorithmic (database schema enables batch processing of dozens of models). The framework is designed for direct incorporation into ALICE, CMS, and JETscape workflows without conceptual modification.
The methodology does not replace traditional $\chi^{2}$/ndf but complements it by revealing physical structure hidden in global statistics. We invite the heavy-ion community to adopt this framework for systematic model evaluation and development of physics-informed hybrid approaches.
Acknowledgements.
We acknowledge the organizers of the 2nd Workshop on Nuclear Physics and High-Energy Physics: Theory, Experiment, and Applications (Kyiv, Ukraine, 2026) for the opportunity to present preliminary results of this work indico2026 .